how do i write a open sentence using absolute value when i have a number line that has these numbers plotted -11 and -7 and then for these two numbers 4 and 12.
Answer by stanbon(75887)

how do i write an open sentence using absolute value when i have a number line that has these numbers plotted -11 and -7
Find the middle of the interval:
It is -9
Write the open sentence describing the distance from -9 as 2:
|x-(-9)| = 2
x+9 = 2 or x+9 = -2
x = -7 or x = -11
-------------------------
And then for these two numbers 4 and 12.
Middle of the interval is 8
Distance from middle is 4
------
Open sentence:
|x-8| = 4
x-8 = 4 or x-8 = -4
x = 12 or x = 4
====================
Cheers,
Stan H.